a father is 4 times as old as his son he will be 3 times as old as his child in 5 years what are their present ages?
Answer by sachi(548)

a father is 4 times as old as his son
let the son is x yrs
so father is 4x yrs
after 5 yrs
the son is x+5 yrs
father is 4x+5 yrs
father will be 3 times as old as his child in 5 years
so 4x+5=3(x+5)=3x+15
or x=15-5=10=sons present age
so father age=4*10=40 yrs